OpenStudy (anonymous):

What is the circumference of a circle with a diameter equaling 7 centimeters in length? Use 3.14 for π. 21.98 cm 31.4 cm 10.99 cm 43.96 cm

OpenStudy (astrophysics):

\[C=2 \pi r\] \[C = 2(3.14)(7)\]

OpenStudy (astrophysics):

Evaluate.

OpenStudy (anonymous):

use circumference = pi d =3.14 * 7=21.98

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome